# Actminer: Applying Causality Tracking and Increment Aligning for Graph-based Threat Hunting

### v1

@@ V-C RQ3: How important are the components we design for assisting threat hunting?
TABLE VIII: The edges and nodes in the query graph, OV and OE denote the number of nodes and edges in the graph before QGP, and V and E denote the number of nodes and edges after processing, respectively.
@@
V-D RQ4: How efficient is Actminer compared with the SOTA in terms of runtime overhead?
We evaluate how efficiently Actminer can detect APT attacks in a timely manner by measuring its running time performance. As shown in Table IX, it can be seen that under the overall time overhead, POIROT consumes 1.69 times more time than Actminer. This suggests that Actminer has better timeliness compared to POIROT. However, in some cases, especially the scenario of E3-Trace case1. Through our further investigation, we have discovered a key cause that led to this discrepancy. Actminer employs a mechanism to offload data instances spanning over 6 hours from memory to the database. While beneficial for optimizing memory management, this storage approach incurs extra overhead when encountering situations that necessitate repeated interactions between the database and memory, resulting in higher time expenditure compared to POIROT.
Furthermore, the presence of an incremental hunting module allows our system to input data in segments for hunting, rather than having to input all the data at once for each hunt. Specifically, enterprises only need to input the newly added data to the system for each batch (e.g., the data from each day), instead of combining it with data from previous days and performing rescanning. At the same time, we have designed a storage mechanism that stores branches in a database. When an edge interacting with that node appears again, the branch is retrieved from the database.
Instead of loading all historical data for each new hunt, only the newly collected information needs to be processed and indexed. As illustrated in the Figure 9, POIROT’s mean memory requirements gradually increase as more data is collected from hosts over time. However, our incremental hunting module ensures that the rate of memory consumption growth is consistently maintained at a steady pace, primarily determined by our storage mechanism.
TABLE IX: Overhead of Actminer and POIROT.
@@ V-D RQ4: How efficient is Actminer compared with the SOTA in terms of runtime overhead?
V-E RQ5: How robust is Actminer in benign dataset?
To comprehensively evaluate the robustness of Actminer, we conducted experiments on a sizeable benign dataset, introduced in Section V. This dataset was collected from multiple users performing typical non-malicious actions such as downloading and uploading files, taking backups, browsing the web, and installing or uninstalling software. Additionally, we included the OpTC dataset, which involves benign activities like website browsing, checking emails, and SSH log-ins. We randomly selected three hosts and tested their benign data collected over a period of one day from the OpTC dataset and five datasets acquired from our own laboratory.
We applied Actminer to all benign datasets and searched for the query graphs extracted from the TC reports. Although these logs were attack-free, they shared many nodes and events with our query graphs, such as critical system files and processes related to email clients and text editing tools. As illustrated in Table X, despite these similarities, Actminer successfully demonstrated robustness by generating zero false alerts throughout the experiment.
TABLE X: Experimental results on different benign datasets.
@@ V-E RQ5: How robust is Actminer in benign dataset?
VI Discussion & Future Work
@@
The accuracy of Actminer. Due to the accuracy of threat hunting relying on the quality of query graphs, we utilized EXTRACTOR to automate the extraction of CTI reports provided by DARPA. We achieved results superior to the SOTA on well-known APT attack datasets (i.e., E3 and E4). However, limited by scarce APT attack samples, we were unable to conduct large-scale experiments and analyses. In practice, accurately extracting query graphs from numerous CTI reports and automatically generating diverse and rational ones can improve threat hunting accuracy in the future.
@@
The interpretability of Actminer. The outputs of Actminer inherently provide interpretability, as analysts can derive attack-related information from the query graphs reports. However, due to the dynamic nature of APT attacks, the query graphs and hunting results are not entirely consistent. Therefore, in the future, we can utilize generative artificial intelligence (e.g., LLM), to automatically transform the attack chains obtained from threat hunting into CTI reports that analysts can clearly understand, aiding in better response.
@@
The robustness of Actminer. The use of graph processing approaches may lead to the loss of fine-grained details. To validate the robustness of our Actminer, we conducted cross-validation by employing query graphs extracted from diverse CTI reports against distinct provenance graphs. The experimental findings consistently demonstrate that our system effectively avoids generating false alarms across all tested origin query graphs. The identified suspicious nodes do not meet the threshold for triggering an alert.
VII RelatedWork
VII-A Provenance Graph-based Threat Hunting
The provenance graph contains causal relationships between system events, and it is a data structure that can be effectively utilized for cyber threat hunting. DeepHunter [39] utilizes Neural Tensor Networks (NTN) to judge the subgraph relationship through graph embedding. Due to the need for multiple-to-multiple traversal comparisons between subgraphs, the efficiency of DeepHunter will decrease as the size of the provenance graph increases. ProvG-Searcher [11] establishes a coarse-grained threat hunting by employing a graph representation learning method on the subgraph, aiming to enhance hunting efficiency. ThreatRaptor [25] achieves extracting structured threat behaviors from OSCTI and automatically synthesizing query statements to search for malicious activities. Unlike the above studies, Actminer, inspired by POIROT [2], tries to align attack graphs extracted from CTI reports with provenance graphs from system logs to find complete attack chain. However, Actminer ia able to significantly reduces false positives, false negatives, and system overhead through an optimized graph alignment algorithm.
VII-B Graph Matching Algorithms
Graph pattern matching refers to the problem of finding similarities between a small query graph and a large target graph, with widespread applications (e.g., database analysis, search engines, software plagiarism detection, and social networks). Based on whether the matching results are completely consistent, algorithms can be divided into two categories: graph isomorphism matching and graph approximate matching. The main idea of graph isomorphism matching algorithms is to iteratively map nodes from the query graph to the target graph one by one. Ullmann et al. [43] proposes a backtracking-based algorithm that enumerates all subgraphs satisfying the matching requirements using a depth-first search approach. However, as the scale of the graph gradually increases, the enumeration range also expands, resulting in low algorithm efficiency. To address this issue, Cordella et al. [44] proposes the VF2 algorithm, which improves algorithm efficiency by incorporating the verification order of query nodes. However, in practical applications, its time complexity is superlinear, and the need for secondary filtering consumes a significant amount of time. Shasha et al. [45] introduces the GraphGrep algorithm, which achieves fast matching by encoding node semantic information. Several other studies [46, 47, 48] use graph mining techniques to find subgraphs from databases and then employ filtering and optimization strategies to prune incorrect nodes. However, due to the diversity and complexity of APT techniques, we cannot always assume that the nodes and edges of the query graph can be fully mapped to the target graph. Graph approximate matching algorithms often rely on heuristic methods to identify important nodes and then gradually expand to neighboring nodes [2, 49, 50, 51]. Tian et al. [49] uses a graph distance model to measure the similarity between graphs. He et al. [52] introduces an index-based algorithm to support subgraph queries and similarity queries. Several other works [53, 51] consider the shape and edge attributes of the query graph. However, the aforementioned research overlooks adversarial knowledge. Milajerdi et al. [2] proposes POIROT, which is similar to our work and utilizes node attributes and information flows between nodes for approximate matching. However, if an attacker intentionally takes a detour to achieve their goal, it may result in missed detections. Furthermore, if the similarity score exceeds the threshold, POIROT stops hunting, and the obtained attack subgraph may not represent the optimal attack behavior. Therefore, we develop a new matching technique to address these challenges.
VII-C Incremental Graph Computation
In real-world scenarios, graphs are typically large in scale and frequently updated over time. When graphs are updated, traditional batch processing methods require starting the computation from scratch, which is extremely time-consuming. In contrast to traditional batch processing algorithms, incremental matching only analyzes and matches the updated portion, utilizing previous matching results to maximize the reduction of redundant computations. Fan et al. proposes the IncSIMMatch [54] algorithm, which effectively reduces redundant computations by creating an index for the pattern graph. They then introduces an incremental computation method IncISO [55], where only the set of nodes within $d$ hops of the updated node $d$ in the data graph needs to be rematched as the affected region when nodes and edges change in the graph. Subsequently, researchers [56, 57, 58, 59, 60, 61, 62] introduce the idea of incremental computation and incorporated query search optimization strategies. For example, Sutanay et al. [57] constructs the pattern graph as a binary tree and decomposes it for storage in tree nodes, with data updates performed by searching leaf nodes. However, storing a large number of indexes consumes memory. Sun et al. [63] proposes an exploration-based approximate matching technique, but if the initially selected node is inappropriate, a large number of useless intermediate result values will be generated. Actminer constructs a tree structure to incrementally update the alignment results and introduces a forgetting rate to maintain stable memory overhead.
VIII Conclusion

### v2

@@ 6.3 RQ3: How important are the components we design for assisting threat hunting?
Table 9: The edges and nodes in the query graph, OV and OE denote the number of nodes and edges in the graph before QGP, and V and E denote the number of nodes and edges after processing, respectively.
@@
6.4 RQ4: How efficient is Actminer compared with the SOTA in terms of runtime overhead?
We evaluate how efficiently Actminer can detect APT attacks in a timely manner by measuring its running time performance. As presented in Table 10, POIROT consumes 1.69 times more time than Actminer under the overall time overhead. This suggests that Actminer exhibits better timeliness compared with POIROT. However, in some cases, especially the scenario of E3 Trace case1. Through our further investigation, we have discovered a key cause that led to this discrepancy. Actminer employs a mechanism to off-load data instances spanning over 6 h from memory to the database. While beneficial for optimizing memory management, this storage approach incurs extra overhead when encountering situations that necessitate repeated interactions between the database and memory, resulting in higher time expenditure compared to POIROT.
Furthermore, the presence of an incremental hunting module allows our system to input data in segments for hunting, rather than having to input all the data at once for each hunt. Specifically, enterprises only need to input the newly added data to the system for each batch (e.g., the data from each day), instead of combining it with data from previous days and performing rescanning. At the same time, we have designed a storage mechanism that stores branches in a database, in which the branch is retrieved from the database when an edge interacting with that node reappears.
Instead of loading all historical data for each new hunt, only the newly collected information needs to be processed and indexed. As illustrated in Figure 11, POIROT’s mean memory requirements gradually increase as more data are collected from hosts over time. However, our incremental hunting module ensures that the rate of memory consumption growth is consistently maintained at a steady pace, primarily determined by our storage mechanism.
Table 10: Overhead of Actminer and POIROT.
@@ 6.4 RQ4: How efficient is Actminer compared with the SOTA in terms of runtime overhead?
6.5 RQ5: How robust is Actminer in benign dataset?
To comprehensively evaluate the robustness of Actminer, we conducted experiments on a sizable benign dataset, introduced in Section 6. This dataset was collected from multiple users performing typical nonmalicious actions, such as downloading and uploading files, taking backups, browsing the Web, and installing or uninstalling software. Additionally, we included the E3, OpTC dataset, which involves benign activities such as website browsing, checking emails, and SSH log-ins.
We randomly selected three hosts and tested their benign data collected over a period of 1 day from the OpTC dataset and five datasets acquired from our own laboratory.
We applied Actminer to all benign datasets and searched for the query graphs extracted from the TC reports. Although these logs were attack-free, they shared many nodes and events with our query graphs, such as critical system files and processes related to email clients and text editing tools. As presented in Table 11, despite these similarities, Actminer successfully demonstrated robustness by generating zero false alerts throughout the experiment.
Table 11: Experimental results on different benign datasets.
@@ 6.5 RQ5: How robust is Actminer in benign dataset?
7 Discussion & Future Work
@@
Accuracy of Actminer. Due to the accuracy of threat hunting relying on the quality of query graphs, we utilized Extractor to automate the extraction of CTI reports provided by DARPA. We achieved results superior to those of SOTA on well-known APT attack datasets (i.e., E3 and E4). However, due to scarce APT attack samples, we were unable to conduct large-scale experiments and analyses. In practice, accurately extracting query graphs from numerous CTI reports and automatically generating diverse and rational ones can improve threat hunting accuracy in the future.
EST of Actminer. We acknowledge that the EST module in Actminer introduces additional computational overhead and may generate more FPs in practical deployments, as demonstrated in our experimental evaluation. However, the fundamental advantage of EST lies in its ability to significantly reduce FNs, thereby minimizing the risk of undetected security threats. In real-world scenarios, security analysts can mitigate the impact of increased FPs through structured alert triage systems that prioritize high-severity incidents and implement automated filtering for common low-risk patterns. The comprehensive coverage of the EST module ensures that subtle attack patterns, which might be missed by traditional detection methods, are captured and flagged for investigation. While this approach requires additional analyst resources for alert processing, the enhanced security posture provided by EST’s reduced FN rate offers substantial long-term benefits.
To address data quality concerns, we recommend implementing robust data collection frameworks, such as Spade gehani2012spade , Kellect chen2022kellect , or eAudit sekar2023eaudit , which provide more complete provenance information compared with legacy tools. Future work should focus on developing intelligent alert correlation mechanisms and automated FP suppression techniques to further optimize the practical applicability of the EST module while maintaining its superior detection capabilities.
@@
Interpretability of Actminer. The outputs of Actminer inherently provide interpretability, as analysts can derive attack-related information from the query graph reports. However, the query graphs and hunting results are not entirely consistent due to the dynamic nature of APT attacks. We can utilize generative artificial intelligence (e.g., large language model) in the future to automatically transform the attack chains obtained from threat hunting into CTI reports that analysts can clearly understand, aiding in better response.
@@
Robustness of Actminer. The use of graph processing approaches may lead to the loss of fine-grained details. To validate the robustness of Actminer, we conducted cross-validation by employing query graphs extracted from diverse CTI reports against distinct provenance graphs. The experimental findings consistently demonstrate that Actminer effectively avoids generating false alarms across all tested origin query graphs. The identified suspicious nodes do not meet the threshold for triggering an alert.
Compression frequency of Actminer. During the provenance graph preparation process, Actminer only retains the most recent timestamp for consecutive temporal events between the same two entities. By retaining only the most recent timestamps, Actminer significantly reduces the need to store redundant data. In consecutive events of the same entity pair, the latest timestamp usually contains the most relevant semantic information. For example, in network threat intelligence, the latest event may indicate the evolution of an attack or a change in strategy. This design choice leads to a critical limitation: Actminer cannot effectively defend against attacks such as ransomware that rely heavily on event frequency patterns for detection. In practical deployment scenarios, such frequency-dependent attacks require the integration of additional specialized tools, such as Bit Defender bitdefender and other complementary security solutions. In future work, it is possible to consider integrating lightweight frequency anomaly detection algorithms as preprocessing steps for provenance graph analysis.
8 Conclusion
@@
9 Acknowledgment
This work is supported in part by the following grants: National Natural Science Foundation of China under Grant No. U22B2028, 62372410, and 62002324. Zhejiang Provincial Natural Science Foundation of China under Grant No. LQ21F020016, and LZ23F020011. The Fundamental Research Funds for the Provincial Universities of Zhejiang under Grant No. RF-A2023009. Wenzhou Key Scientific and Technological Projects under Grant No. ZG2024007. Wenzhou Basic Scientific Research Projects under Grant No. G2024033. Open Fund Project of Wenzhou Cyber Security Detection and Protection Technology Research Center under Grant No. WZKF2023001, and WZKF2023002. "Pioneer" and "Leading Goose" R&D Program of Zhejiang under Grant No. 2025C01082, and 2025C01013.